Let the functional f on R² be defined by f(x)= 4x-3y.Regard R² as a subspace of R³ given by z=0 determine all linear extension of f(x) from R² to R³


1
Expert's answer
2021-03-01T07:21:07-0500

The of functional "f(x)" on "R^2" :

"||f||_{R^2}=\sqrt{4^2+3^2}=5"

The norm of linear extension:

"||f_0||_{R^3}=||f||_{R^2}"


So, linear extension is:

"f_0=a_1x+a_2y+a_3z"

"\sqrt{a_1^2+a_2^2+a_3^2}=5"